. A tailor has the following material available: 16 sq. yd. cotton, 11 sq. yd. silk and 15 sq. yd. wool. A suit requires 2 sq. yd. cotton, 1 sq. yd. silk and 1 sq. yd. wool. A gown requires 1 sq. yd. cotton, 2 sq. td. Silk and 3 sq. yd. wool. If a suit sells for Rs. 300/- and a gown for Rs. 500/-, how many of each garment should the tailor mak to obtain the maximum amount of profit?


1
Expert's answer
2021-09-20T04:52:14-0400

Let "x=" the number of suits, "y=" the number of gowns.

Maximize "Z=300x+500y"

subject to


"2x+y\leq16"

"x+2y\leq11"

"x+3y\leq 15"

"x\geq0, y\geq0"

"OA: x=0, 0\leq y\leq5, Z=500y"

"AB: y=5-\dfrac{1}{3}x, 0\leq x\leq3, Z=\dfrac{400}{3}x+2500"

"BC: y=\dfrac{11}{2}-\dfrac{1}{2}x, 3\leq x\leq7, Z=50x+2750"

"CD: y=16-2x, 7\leq x\leq8, Z=-700x+8000"


"OD: y=0, 0\leq x\leq8, Z=300x"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Extreme & Lines\ through & Objective \ function \\
point & exreme\ point & value\\ \hline
 O (0,0) & x=0, & Z=500(0) \\ 
& y=0 &=0\\
 \hdashline
 A (0,5) & x=0, & Z=500(5)\\
 & y=5-\dfrac{1}{3}x & =2500\\
 \hdashline
B (3,4) & y=5-x/3, & Z=50(3)+2750\\
 & y=11/2-x/2 & =2900\\
 \hdashline
 C (7,2) & y=11/2-x/2, & Z=-700(7)+8000\\
 & y=8-2x & =3100\\
 \hdashline
 D (8,0) & y=8-2x & Z=300(8)\\
 & y=0 & =2400\\
 \hdashline
\end{array}"

The maximum value of the objective function "Z=3100"  occurs at the extreme point "(7, 2)."

Hence, the optimal solution to the given LP problem is :

 "x=7,y=2"  and "max Z=3100."